# Network Alignment with Transferable Graph Autoencoders

### v2

@@ E.4 T-GAE Parameters
To select the hyperparameters of our model we utilize a validation graph (JAZZ (Rossi & Ahmed, 2015)), which is not included in our comparisons to ensure unbiased comparisons. Note that all parameters of T-GAE are fixed for all datasets that are reported. The model is not fine-tuned on every network that is used in testing.
The parameters of T-GAE include:
  • We train our model Adam optimizer for 100 epochs and initial learning rate equal to $0.001$ for graph matching experiment. To have a fair comparision with WAlign (Gao et al., 2021), we train 20 epochs for the subgraph matching tasks.
  • The hidden dimension of the GNN modules is equal to $6$ and that of the output layer (encoder) is equal to 4 for graph matching experiments. For subgraph matching, the hidden dimensions are set to be 512, the same to the optimal value for WAlign (Gao et al., 2021).
  • The activation function for each GNN layer is ReLU.
  • We use an 1-layer decoder followed by the sigmoid function.
  • To conduct the experiments in Sections G.0.1, G.3 we augment the dataset with $25$ perturbed graphs for each perturbation level.
Appendix F Efficiency and Run-time Comparison
We analyze the efficiency of the proposed graph matching framework by comparing its running time with the competing algorithms. The GNN-based methods achieve at most $\times$50 less running time on both graph matching and subgraph matching tasks. The proposed training objective (8) scales well from networks with 1,000 nodes (Zhang & Tong, 2016) to denser networks with $\times$10 nodes, with minor running time increase, compared to other node embedding techniques.
We also noticed that compared to the other GNN based framework WAlign (Gao et al., 2021), T-GAE takes more time to train on the larger scale networks as illustrated in Figure 6. This is because the deep GNN is trained by the structure of the network. The training takes more time as the netowrk scales up and the number of nodes and edges grow dramatically, whereas the training labels for WAlign come from the input feature, which does not scale up with as the size of the training graphs increase. This could be one of the limitations to the proposed training paradigm and that is why we leverage transfer learning to efficiently match large scale graphs, as illustrated in Table3.
Algorithm
Celegans
Arenas
Douban
Cora
Dblp
Coauthor CS
T-GAE
0.0980.098{0.098}0.098
0.2410.241{0.241}0.241
1.9341.934{1.934}1.934
0.8420.842{0.842}0.842
173.836173.836{173.836}173.836
193.289193.289{193.289}193.289
GAE
0.0940.094{0.094}0.094
0.2520.252{0.252}0.252
1.9491.949{1.949}1.949
0.8570.857{0.857}0.857
174.410174.410{174.410}174.410
194.062194.062{194.062}194.062
WAlign
0.0780.078{0.078}0.078
0.2050.205{0.205}0.205
1.8001.800{1.800}1.800
0.7660.766{0.766}0.766
169.694169.694{169.694}169.694
189.032189.032{189.032}189.032
Spectrul
5.7125.712{5.712}5.712
2.8192.819{2.819}2.819
60.29860.298{60.298}60.298
54.77054.770{54.770}54.770
> 10 hours
> 10 hours
GraphWave
8.3088.308{8.308}8.308
32.99432.994{32.994}32.994
281.629281.629{281.629}281.629
131.230131.230{131.230}131.230
5470.7245470.724{5470.724}5470.724
6291.3686291.368{6291.368}6291.368
ConeAlign
1.3331.333{1.333}1.333
3.5003.500{3.500}3.500
31.95531.955{31.955}31.955
13.79913.799{13.799}13.799
887.090887.090{887.090}887.090
1099.1451099.145{1099.145}1099.145
S-GWL
27.84427.844{27.844}27.844
37.44337.443{37.443}37.443
3394.5223394.522{3394.522}3394.522
311.201311.201{311.201}311.201
> 10 hours
> 10 hours
FINAL
0.0300.030{0.030}0.030
0.0810.081{0.081}0.081
1.0071.007{1.007}1.007
0.4980.498{0.498}0.498
86.78886.788{86.788}86.788
118.065118.065{118.065}118.065
Appendix G Uniform Perturbation Model Supplementary Results
G.0.1 Perturbed Training
In the previous experiment, T-GAE is trained with a family of original graphs and tested to match perturbed versions of a larger family of graphs. T-GAE exhibited more robust performance compared to the baseline methods, however, its matching accuracy dropped significantly as the perturbation level of testing data increased. To tackle this problem we follow a self-supervised learning approach and train T-GAE with a family of real graphs and perturbations of them. We train according to (8), which aims to produce similar node embeddings for both the original graphs and perturbed versions of them. The data augmentation process follows the previously explained perturbation models. Similar to the previous experiment we train over the four small datasets and execute over all datasets. Note that training and testing is performed with different perturbations of the original graphs. Table 3 reports the testing results of the best T-GAE when training is performed with graph perturbations.
Table 3: Performance Comparison of T-GAE when trained with/without perturbation.
Algorithm
Celegans
Arenas
Douban
Cora
Dblp
Coauthor CS
0% perturbation
T-GAE
${89.5\pm 1.3}$
${88.4\pm 0.5}$
${90.3\pm 0.4}$
${87.4\pm 0.4}$
${85.6\pm 0.1}$
${97.6\pm 0.1}$
T-GAE with pert.
$\mathbf{89.7\pm 1.5}$
$\mathbf{88.6\pm 0.6}$
$\mathbf{90.1\pm 0.4}$
$\mathbf{87.4\pm 0.5}$
$\mathbf{85.7\pm 0.2}$
$\mathbf{97.7\pm 0.1}$
1% perturbation
T-GAE
$\mathbf{84.1\pm 1.1}$
${84.8\pm 0.6}$
${84.9\pm 0.6}$
${82.9\pm 0.5}$
${79.1\pm 0.4}$
${86.5\pm 0.8}$
T-GAE with pert.
${83.4\pm 1.6}$
$\mathbf{85.6\pm 0.5}$
$\mathbf{85.2\pm 0.6}$
$\mathbf{83.2\pm 0.9}$
$\mathbf{79.7\pm 0.4}$
$\mathbf{87.1\pm 1.0}$
5% perturbation
T-GAE
${50.8\pm 3.3}$
${47.1\pm 5.6}$
${57.9\pm 6.1}$
$\mathbf{58.2\pm 2.0}$
${40.8\pm 2.1}$
${26.9\pm 5.4}$
T-GAE with pert.
$\mathbf{52.3\pm 5.4}$
$\mathbf{62.6\pm 2.2}$
$\mathbf{58.5\pm 5.0}$
${57.4\pm 2.7}$
$\mathbf{43.6\pm 3.7}$
$\mathbf{30.4\pm 7.4}$
We observe that incorporating graph perturbations in the training process significantly helps in high perturbation levels and benefits the robustness of the proposed method. On the other hand, when testing on low levels of perturbations, using the original graph or perturbations to train the T-GAE does not lead to significant changes. In particular, at 5% testing perturbation, T-GAE archives 15.5% increase in the Arenas dataset, whereas at 0% and 1% testing perturbation the increase is 0.2% and 0.8% respectively.
[…]
We present the graph matching accuracy of the proposed T-GAE method using different message passing mechanisms as the backbone of the encoder in Table 4. The experiment setup is the same as in Table 3, we train on the 4 smaller datasets (Celegans, Arenas, Douban, Cora) and test on 10 randomly generated perturbations of all the 6 datasets. We observe that the proposed framework work well with different types of message passing GNNs in graph matching tasks and is able to produce robust node embedding at small perturbation levels.
[…]
Algorithm
Celegans
Arenas
Douban
Cora
Dblp
Coauthor CS
$0\%$
[…]
${89.5\pm 1.3}$ ${88.4\pm 0.5}$ ${90.1\pm 0.4}$ ${87.1\pm 0.6}$
${85.6\pm 0.1}$
${97.6\pm 0.1}$
T-GAE (GCN) ${87.8\pm 1.1}$ ${88.3\pm 0.4}$ ${89.9\pm 0.2}$ ${87.4\pm 0.4}$ ${85.8\pm 0.1}$ ${97.6\pm 0.1}$
[…]
T-GAE (GCN) ${83.4\pm 2.3}$ ${84.8\pm 0.6}$ ${84.1\pm 1.2}$ ${82.5\pm 0.8}$ ${76.0\pm 1.1}$ ${84.3\pm 0.8}$
[…]
${21.7\pm 5.2}$
T-GAE (GCN)
${37.5\pm 9.7}$
${45.0\pm 8.2}$
${49.5\pm 4.2}$
${50.9\pm 2.8}$
${31.8\pm 2.8}$
${23.0\pm 4.1}$
G.2 More Baseline Results
Table 5: Graph matching accuracy on 10 randomly perturbed samples under different levels of edge editing for VGAE, LINE and DeepWalk
Dataset
VGAE
LINE
DeepWalk
$0\%$ Celegans ${0.3\pm 0.1}$ ${1.0\pm 0.5}$ ${1.8\pm 0.6}$
Arenas
${0.1\pm 0.1}$
${0.2\pm 0.1}$
${0.3\pm 0.2}$
Douban
${0.0\pm 0.0}$
${0.0\pm 0.0}$
${0.1\pm 0.0}$
Cora
${0.1\pm 0.0}$
${0.0\pm 0.0}$
${0.1\pm 0.0}$
$1\%$ Celegans ${0.3\pm 0.1}$ ${1.0\pm 0.4}$ ${1.2\pm 0.5}$
Arenas
${0.1\pm 0.1}$
${0.1\pm 0.1}$
${0.3\pm 0.1}$
Douban
${0.0\pm 0.0}$
${0.0\pm 0.0}$
${0.1\pm 0.0}$
Cora
${0.1\pm 0.1}$
${0.1\pm 0.0}$
${0.2\pm 0.1}$
$5\%$ Celegans ${0.6\pm 0.3}$ ${0.9\pm 0.3}$ ${1.0\pm 0.3}$
Arenas
${0.2\pm 0.1}$
${0.2\pm 0.2}$
${0.2\pm 0.1}$
Douban
${0.0\pm 0.0}$
${0.0\pm 0.0}$
${0.0\pm 0.0}$
Cora
${0.1\pm 0.0}$
${0.1\pm 0.0}$
${0.1\pm 0.0}$
Hit rate
GraphWave
Hit@1
0.090.09{0.09}0.09
Hit@5
0.360.36{0.36}0.36
Hit@10
0.810.81{0.81}0.81
Hit@50
4.744.74{4.74}4.74
Hit@100
9.129.12{9.12}9.12
G.3 Specific Perturbed Training
The last question that we will try to answer with our experiments about Uniform Perturbation model is whether training and testing on the same graph compares to training with a family of graphs or a combination of a family and perturbed graphs, that was examined in the previous subsection. In this subsection we test the performance of T-GAE on Celegans, Arenas, Douban, and Cora, on 4 different settings: i) when training on each graph separately according to (6); ii) when training on all 4 graphs according to (7); iii) when training on the perturbations of only one graph according to (7), but $\mathbb{S}$ being defined over a family of graph perturbations; iv when training on all 4 graphs plus perturbed versions of them, according to (8). The results for the different levels of perturbations are presented in Table 7.
Table 7: Performance of T-GAE when trained on a family of graphs or on a specific graph on uniform perturbation model
Perturbation
Celegans
Arenas
Douban
Cora
0% perturbation
T-GAE specific (6)
${89.3\pm 1.4}$
${88.3\pm 0.4}$
${90.2\pm 0.5}$
${87.5\pm 0.4}$
T-GAE (7)
${89.5\pm 1.3}$
${88.4\pm 0.5}$
${90.1\pm 0.4}$
${87.4\pm 0.4}$
T-GAE specific with pert.
$\mathbf{90.4\pm 0.8}$
${88.5\pm 0.4}$
$\mathbf{90.2\pm 0.3}$
$\mathbf{{87.6\pm 0.5}}$
T-GAE with pert. (8)
${89.7\pm 1.5}$
$\mathbf{88.6\pm 0.6}$
${90.1\pm 0.4}$
${87.4\pm 0.5}$
1% perturbation
T-GAE specific (6)
${81.3\pm 1.7}$
${85.4\pm 0.5}$
${85.1\pm 1.1}$
${83.1\pm 1.0}$
T-GAE (7)
$\mathbf{84.1\pm 1.1}$
${84.8\pm 0.6}$
${84.9\pm 0.6}$
${82.9\pm 0.5}$
T-GAE specific with pert.
${83.1\pm 1.4}$
${85.3\pm 0.9}$
$\mathbf{85.3\pm 0.7}$
$\mathbf{83.5\pm 0.6}$
T-GAE with pert. (8)
${83.4\pm 1.6}$
$\mathbf{85.6\pm 0.5}$
${85.2\pm 0.6}$
${83.2\pm 0.9}$
5% perturbation
T-GAE specific (6)
${48.9\pm 9.5}$
${61.5\pm 5.4}$
${58.0\pm 3.9}$
${58.8\pm 2.7}$
T-GAE (7)
${50.8\pm 3.3}$
${47.1\pm 5.6}$
${57.9\pm 6.1}$
${58.2\pm 2.0}$
T-GAE specific with pert.
$\mathbf{53.5\pm 11.9}$
${54.3\pm 6.4}$
${57.3\pm 4.6}$
$\mathbf{58.9\pm 2.4}$
T-GAE with pert. (8)
${52.3\pm 5.4}$
$\mathbf{62.6\pm 2.2}$
$\mathbf{58.5\pm 5.0}$
${57.4\pm 2.7}$
Table 7 shows when testing on low perturbation or no perturbations, all the models work statistically the same. The only exception is in the Celegans dataset where training on a family of graphs seems to improve the matching performance by $2.8\%$ compared to training just on a single graph. For $5\%$ perturbation level, on the other hand, either T-GAE specific with pert. or T-GAE with pert. yield the best results, which implies that a self-supervised approach with perturbations benefits performance and robustness. We also observe that T-GAE with pert., i.e., training according to (8) is in general the most consistent approach.
Appendix H Degree Perturbation Model Results
@@ H.1 Transferability Analysis
We test the performance of T-GAE for large-scale network alignment on the degree perturbation model, as described in Section 5.3. We adopt the same setting as in Section 5.3 to train the T-GAE according to (7) on small-size networks, i.e., Celegans, Arena, Douban, and Cora, and conduct transfer learning experiments on the larger graphs, i.e., Dblp, and Coauthor CS. The trained T-GAE is used to generate node embedding for the graphs, and 1 computes the assignment matrix. The results presented in Table 8 are based on the average and standard deviation of the matching accuracy over 10 randomly generated perturbed samples.
Table 8: Graph Matching Accuracy for degree perturbation model
Algorithm
Celegans
Arenas
Douban
Cora
Dblp
Coauthor CS
$0\%$ perturbation Spectral ${88.1\pm 1.1}$ ${97.5\pm 0.5}$ ${90.0\pm 0.4}$ ${84.7\pm 0.5}$
${84.6\pm 0.1}$
$\mathbf{97.6\pm 0.1}$
NetSimile ${71.8\pm 1.0}$ ${94.9\pm 0.4}$
${46.9\pm 0.5}$
${73.0\pm 0.7}$
${63.7\pm 0.2}$
${90.1\pm 0.1}$
FINAL $\mathbf{92.7\pm 1.9}$ ${97.5\pm 0.5}$ ${90.0\pm 0.4}$ ${87.1\pm 0.5}$ $\mathbf{85.6\pm 0.1}$ $\mathbf{97.6\pm 0.1}$
[…]
${89.3\pm 1.2}$ $\mathbf{97.9\pm 0.5}$ $\mathbf{90.2\pm 0.4}$ ${87.5\pm 0.5}$ $\mathbf{85.6\pm 0.1}$ $\mathbf{97.6\pm 0.1}$
T-GAE (GCN) ${89.7\pm 1.4}$ ${97.5\pm 0.3}$ ${90.1\pm 0.3}$ ${87.4\pm 0.4}$ $\mathbf{85.6\pm 0.1}$ $\mathbf{97.6\pm 0.1}$
T-GAE (GIN) ${89.6\pm 1.0}$ ${97.6\pm 0.4}$ ${90.1\pm 0.3}$ $\mathbf{87.6\pm 0.4}$ ${85.4\pm 0.1}$ ${97.5\pm 0.1}$
$1\%$ perturbation Spectral ${21.8\pm 16.9}$ ${66.0\pm 21.6}$
${9.8\pm 13.5}$
${25.3\pm 13.4}$ ${3.4\pm 0.8}$
${8.6\pm 4.6}$
NetSimile ${63.1\pm 1.4}$
${90.9\pm 0.7}$
${37.6\pm 0.6}$
${64.4\pm 1.2}$ ${52.2\pm 0.5}$
${76.8\pm 0.6}$
FINAL
${62.5\pm 4.1}$
${56.7\pm 2.7}$ ${35.0\pm 1.1}$ ${32.0\pm 2.0}$
${24.5\pm 0.6}$
${19.1\pm 0.5}$
[…]
T-GAE (GCN) $\mathbf{68.9\pm 5.2}$ ${86.7\pm 2.9}$ ${82.2\pm 1.4}$ ${78.5\pm 1.3}$ ${62.3\pm 1.6}$ ${83.6\pm 1.0}$
T-GAE (GIN) ${49.6\pm 16.2}$ $\mathbf{92.7\pm 1.2}$ $\mathbf{83.7\pm 2.2}$ $\mathbf{81.0\pm 2.5}$ $\mathbf{75.1\pm 0.6}$ $\mathbf{88.3\pm 0.4}$
$5\%$ perturbation Spectral
${14.4\pm 3.9}$
${28.3\pm 10.0}$
${1.4\pm 0.7}$ ${3.5\pm 1.9}$ ${0.3\pm 0.1}$
${0.5\pm 0.2}$
NetSimile
${12.6\pm 1.4}$
${22.4\pm 1.0}$
${10.3\pm 0.4}$
${32.0\pm 0.7}$
${18.4\pm 0.1}$
${20.2\pm 0.3}$
FINAL $\mathbf{24.3\pm 2.3}$ ${20.5\pm 1.6}$
${7.5\pm 0.4}$
${8,6\pm 0.7}$
${4.7\pm 0.2}$
${3.5\pm 1.3}$
[…]
${10.8\pm 0.4}$
T-GAE (GCN)
${9.9\pm 2.6}$
${28.3\pm 1.3}$
${30.7\pm 1.4}$
${47.0\pm 1.7}$
${22.2\pm 0.7}$
${13.6\pm 0.5}$
T-GAE (GIN)
${13.0\pm 1.8}$
$\mathbf{42.9\pm 2.1}$
${28.9\pm 1.4}$
$\mathbf{52.4\pm 3.3}$
$\mathbf{35.7\pm 0.8}$
$\mathbf{23.6\pm 0.6}$
When testing on 0 perturbation level most algorithms are able to discover the node correspondences, as in the uniform perturbation model. The benefit of processing the NetSimile embeddings with GNNs is still significant in this perturbation model as we observe up to $53\%$ performance increase. When testing on perturbed graphs, our proposed T-GAE consistently outperforms all the competing baselines, while being robust and efficient when performing network alignment under different perturbation models. Finally, on large-scale networks, T-GAE is able to achieve very high levels of matching accuracy for both Dblp and Coauthor CS, for $p=0\%,~{}1\%$.
H.2 Perturbed Training
We train the T-GAE with a family of original graphs and test to match perturbed versions of a larger family of graphs. We notice that although being more robust to perturbations than the baseline methods, the matching accuracy of T-GAE dropped as the perturbation level of testing data increased, especially on the degree perturbation model. To tackle this problem we follow a self-supervised learning approach and train T-GAE with a family of real graphs and perturbations of them. We train according to (8), which encourages the model to produce similar node embeddings for the original graphs and the perturbed versions of them. We augment the dataset as illustrated in 5, adding and deleting edges, according to the degree Model perturbation probability presented in Section 5.3. Training and testing are conducted on different randomly generated perturbed samples. Table 9 reports the testing results of T-GAE when trained on the best perturbation levels.
Table 9: Performance Comparison of T-GAE when trained with/without perturbation on degree perturbation model
Algorithm
Celegans
Arenas
Douban
Cora
Dblp
Coauthor CS
0% perturbation
T-GAE
${89.7\pm 1.4}$
${88.2\pm 0.4}$
$\mathbf{90.2\pm 0.4}$
${87.6\pm 0.4}$
${85.6\pm 0.1}$
${97.6\pm 0.1}$
T-GAE with pert.
$\mathbf{90.4\pm 1.2}$
$\mathbf{88.4\pm 0.4}$
${90.1\pm 0.3}$
$\mathbf{87.6\pm 0.4}$
$\mathbf{85.7\pm 0.1}$
$\mathbf{97.6\pm 0.0}$
1% perturbation
T-GAE
$\mathbf{68.9\pm 5.2}$
${86.3\pm 0.7}$
${83.7\pm 2.2}$
${81.0\pm 2.5}$
$\mathbf{75.1\pm 0.6}$
${88.3\pm 0.4}$
T-GAE with pert.
${67.2\pm 10.2}$
$\mathbf{87.3\pm 1.0}$
$\mathbf{85.3\pm 1.0}$
$\mathbf{82.7\pm 0.8}$
${74.1\pm 0.7}$
$\mathbf{89.7\pm 0.7}$
5% perturbation
T-GAE
${14.0\pm 2.9}$
${42.9\pm 2.1}$
${30.7\pm 1.4}$
${52.4\pm 3.3}$
${35.7\pm 0.8}$
${23.6\pm 0.6}$
T-GAE with pert.
$\mathbf{17.7\pm 2.0}$
$\mathbf{55.7\pm 3.2}$
$\mathbf{47.1\pm 2.2}$
$\mathbf{57.6\pm 1.1}$
$\mathbf{39.4\pm 0.7}$
$\mathbf{30.2\pm 0.7}$
The results show that introducing training perturbations significantly improves the graph matching accuracy in high testing perturbation levels. When testing on low perturbation levels, however, the use of perturbations to train the T-GAE does not lead to significant changes in this perturbation model. At 5% testing perturbation, training with perturbations offers a 16.4% accuracy increase in the Douban dataset and 12.8% accuracy improvement in the Arenas Dataset, whereas at 1% testing perturbation, the increase is 1% and 1.6% respectively. And at 0 testing perturbation, the performance on these two datasets is similar.
H.3 Specific Perturbed Training
The last experiment we conduct on the degree perturbation model is to compare the performance of T-GAE on Celegans, Arenas, Douban, and Cora, on 4 different settings: i) training is performed on a specific graph according to (6); ii) training on the family of all 4 graphs based on (7); iii) training on the perturbations of one specific graph according to (7), with $\mathbb{S}$ being defined over a family of graph perturbations; iv) training on the perturbations of all 4 graphs according to (8). The results for different levels of perturbations are presented in Table 7.
Table 10: Performance of T-GAE when trained on a family of graphs or on a specific graph on degree perturbation model
Perturbation
Celegans
Arenas
Douban
Cora
0% perturbation
T-GAE specific (6)
${88.9\pm 1.1}$
${88.4\pm 0.5}$
${90.0\pm 0.3}$
${87.5\pm 0.4}$
T-GAE (7)
${89.7\pm 1.4}$
${88.2\pm 0.4}$
$\mathbf{90.2\pm 0.4}$
$\mathbf{87.6\pm 0.4}$
T-GAE specific with pert.
${89.4\pm 1.0}$
${88.3\pm 0.4}$
${90.1\pm 0.4}$
${{87.5\pm 0.6}}$
T-GAE with pert. (8)
$\mathbf{90.4\pm 1.2}$
$\mathbf{88.4\pm 0.4}$
${90.1\pm 0.3}$
$\mathbf{87.6\pm 0.4}$
1% perturbation
T-GAE specific (6)
$\mathbf{74.8\pm 2.8}$
${85.8\pm 1.5}$
${85.1\pm 0.7}$
${82.3\pm 1.3}$
T-GAE (7)
${68.9\pm 5.2}$
${86.3\pm 0.7}$
${83.7\pm 2.2}$
${81.0\pm 2.5}$
T-GAE specific with pert.
${70.7\pm 4.6}$
${87.0\pm 0.7}$
${84.1\pm 0.8}$
${81.6\pm 1.5}$
T-GAE with pert. (8)
${67.2\pm 10.2}$
$\mathbf{87.3\pm 1.0}$
$\mathbf{85.3\pm 1.0}$
$\mathbf{82.7\pm 0.8}$
5% perturbation
T-GAE specific (6)
$\mathbf{30.6\pm 5.0}$
${44.2\pm 1.2}$
${41.3\pm 1.6}$
${56.0\pm 1.4}$
T-GAE (7)
${14.0\pm 2.9}$
${42.9\pm 2.1}$
${30.7\pm 1.4}$
${52.4\pm 3.3}$
T-GAE specific with pert.
${27.3\pm 4.8}$
${46.4\pm 3.8}$
${42.0\pm 2.1}$
${56.3\pm 1.8}$
T-GAE with pert. (8)
${17.7\pm 2.0}$
$\mathbf{55.7\pm 3.2}$
$\mathbf{47.1\pm 2.2}$
$\mathbf{57.6\pm 1.1}$
From Table 10 we observe that all the models work statistically the same when testing on zero perturbation. When testing on low or high perturbation levels, training with perturbation achieves the best accuracy in most datasets. The Celegans dataset is an exception and achieves a better matching accuracy by $5.9\%$ and $3.3\%$ when trained on a specific graph, respectively, compared to training on a family of graphs or perturbed adjacencies. Overall, we see that training with perturbations helps improve the matching accuracy and the robustness of the proposed method. T-GAE with pert. or T-GAE specific with pert.. In other words, training according to (8) is in general produces the most consistent and accurate results compared to the settings where we train the model on the original graph without perturbation.
From the additional perturbation model, we conclude that our proposed T-GAE framework is able to transfer and perform network alignment on large and small-size graphs and is robust to achieve consistent accuracy under different perturbation models and perturbation levels.

### v3

@@ Appendix H Degree Perturbation Model Results
Spectral Netsimile GraphWave FINAL S-GWL ConeAlign WAlign GAE T-GAE
$0\%$ perturbation Celegans ${87.8\pm 1.5}$ ${72.7\pm 0.9}$ ${65.3\pm 1.7}$ ${92.2\pm 1.2}$ $\mathbf{93.0\pm 1.5}$ ${66.6\pm 1.2}$ ${88.4\pm 1.6}$ ${90.9\pm 2.6}$ ${91.0\pm 1.1}$
Arenas ${97.7\pm 0.4}$ ${94.7\pm 0.3}$ ${81.7\pm 0.7}$ ${97.5\pm 0.3}$ ${97.5\pm 0.3}$ ${87.8\pm 0.6}$ ${97.4\pm 0.5}$ ${97.6\pm 0.4}$ $\mathbf{97.8\pm 0.4}$
Douban ${89.9\pm 0.4}$ ${46.4\pm 0.4}$ ${17.5\pm 0.2}$ ${89.9\pm 0.3}$ $\mathbf{90.1\pm 0.3}$ ${68.1\pm 0.4}$ ${90.0\pm 0.4}$ ${89.5\pm 0.4}$ $\mathbf{90.1\pm 0.3}$
Cora ${85.0\pm 0.4}$ ${73.7\pm 0.4}$ ${8.3\pm 0.4}$ ${87.5\pm 0.7}$ ${87.3\pm 0.7}$ ${38.5\pm 0.7}$ ${87.2\pm 0.4}$ ${87.1\pm 0.8}$ $\mathbf{87.5\pm 0.4}$
Dblp ${84.5\pm 0.1}$ ${63.7\pm 0.2}$ doesn’t scale $\mathbf{85.6\pm 0.2}$ > 48 hours ${44.3\pm 0.6}$ ${85.6\pm 0.2}$ ${85.2\pm 0.3}$ $\mathbf{85.6\pm 0.2}$
Coauthor CS ${97.5\pm 0.1}$ ${90.9\pm 0.1}$ doesn’t scale $\mathbf{97.6\pm 0.1}$ > 48 hours ${75.8\pm 0.5}$ ${97.5\pm 0.2}$ ${97.6\pm 0.3}$ $\mathbf{97.6\pm 0.1}$
$1\%$ perturbation Celegans ${21.8\pm 16.9}$ ${63.1\pm 1.4}$ ${4.9\pm 1.1}$ ${62.5\pm 4.1}$ $\mathbf{73.1\pm 11.7}$ ${61.4\pm 3.2}$ ${70.7\pm 4.4}$ ${7.8\pm 2.5}$ ${63.4\pm 6.1}$
Arenas ${66.0\pm 21.6}$ ${90.9\pm 0.7}$ ${5.8\pm 2.6}$ ${56.7\pm 2.7}$ ${92.3\pm 1.2}$ ${86.1\pm 0.6}$ ${96.0\pm 0.5}$ ${0.9\pm 0.5}$ $\mathbf{96.7\pm 0.6}$
Douban ${9.8\pm 13.5}$ ${37.6\pm 0.6}$ ${0.6\pm 0.3}$ ${35.0\pm 1.1}$ ${69.9\pm 1.2}$ ${59.6\pm 2.2}$ ${81.3\pm 1.9}$ ${0.4\pm 0.0}$ $\mathbf{83.7\pm 2.2}$
Cora ${25.3\pm 13.4}$ ${64.4\pm 1.2}$ ${0.1\pm 0.0}$ ${32.0\pm 2.0}$ ${31.1\pm 4.4}$ ${30.6\pm 2.7}$ ${74.6\pm 0.7}$ ${28.2\pm 0.3}$ $\mathbf{81.0\pm 2.5}$
Dblp ${3.4\pm 0.8}$ ${52.2\pm 0.5}$ doesn’t scale ${24.5\pm 0.6}$ > 48 hours ${21.1\pm 1.9}$ ${67.4\pm 1.1}$ ${10.5\pm 0.6}$ $\mathbf{77.1\pm 0.6}$
Coauthor CS ${8.6\pm 4.6}$ ${76.8\pm 0.6}$ doesn’t scale ${19.1\pm 0.5}$ > 48 hours ${64.0\pm 2.0}$ ${88.9\pm 0.8}$ ${3.8\pm 0.1}$ $\mathbf{94.0\pm 0.4}$
$5\%$ perturbation Celegans ${14.4\pm 3.9}$ ${12.6\pm 1.4}$ ${2.6\pm 0.6}$ ${24.3\pm 2.3}$ ${39.1\pm 21.1}$ $\mathbf{60.3\pm 1.9}$ ${6.8\pm 0.9}$ ${2.0\pm 0.7}$ ${18.3\pm 1.4}$
Arenas ${28.3\pm 10.0}$ ${22.4\pm 1.0}$ ${0.5\pm 0.4}$ ${20.5\pm 1.6}$ $\mathbf{92.2\pm 0.9}$ ${82.2\pm 1.2}$ ${6.9\pm 1.1}$ ${0.0\pm 0.0}$ ${60.1\pm 1.9}$
Douban ${1.4\pm 0.7}$ ${10.3\pm 0.4}$ ${0.3\pm 0.0}$ ${7.5\pm 0.4}$ $\mathbf{56.1\pm 12.3}$ ${48.2\pm 2.2}$ ${5.2\pm 0.9}$ ${0.1\pm 0.0}$ ${29.2\pm 2.3}$
Cora ${3.5\pm 1.9}$ ${32.0\pm 0.7}$ ${0.1\pm 0.0}$ ${8.6\pm 0.7}$ ${15.5\pm 5.5}$ ${24.7\pm 3.0}$ ${24.2\pm 1.4}$ ${4.0\pm 0.5}$ $\mathbf{48.5\pm 1.2}$
Dblp ${0.3\pm 0.1}$ ${18.4\pm 0.1}$ doesn’t scale ${4.7\pm 0.2}$ > 48 hours ${10.0\pm 0.8}$ ${14.9\pm 0.2}$ ${5.0\pm 0.3}$ $\mathbf{36.4\pm 1.2}$
Coauthor CS ${0.5\pm 0.2}$ ${20.2\pm 0.3}$ doesn’t scale ${3.5\pm 1.3}$ > 48 hours $\mathbf{47.7\pm 2.7}$ ${4.6\pm 0.1}$ ${0.5\pm 0.1}$ ${35.3\pm 0.5}$
@@
We test the performance of T-GAE for large-scale network alignment on the degree perturbation model, as described in Section 5.3. We adopt the same setting as in Section 5.3 to train the T-GAE according to (7) on small-size networks, i.e., Celegans, Arena, Douban, and Cora, and conduct transfer learning experiments on the larger graphs, i.e., Dblp, and Coauthor CS. The trained T-GAE is used to generate node embedding for the graphs, and Algorithm 1 computes the assignment matrix. The results presented in Table 5 are based on the average and standard deviation of the matching accuracy over 10 randomly generated perturbed samples.
We observe that the benefit of processing the NetSimile embeddings with GNNs is still significant in this perturbation model as we observe up to $46\%$ performance increase at the presence of perturbation. When testing on perturbed graphs at $1\%$ level of edge removal, our proposed T-GAE consistently outperforms all the competing baselines, while being robust and efficient when performing network alignment under different perturbation models. Especially, on large-scale networks, T-GAE is able to achieve very high levels of matching accuracy for both Dblp and Coauthor CS, for $p=0\%,~{}1\%$.
It is also noticed that when we introduce $5\%$ perturbation, the performance of all competing methods dropped compared to the results in Table 3. This is because the Uniform model is agnostic to the graph structure. In the Degree perturbation model however, an edge with higher degree is more likely to be removed, so it changes the structure of the graph more drastically, which makes the task of graph matching more challenging.
The benefit of our proposed T-GAE framework in improving the expressiveness of GNN still stands out if we compare the accuracy with WAlign and GAE. It also consistently achieves the best result among all baseline methods that are salable.